# Enhancing Startup Success Predictions in Venture Capital: A GraphRAG Augmented Multivariate Time Series Method

### v2

@@ Introduction
The algorithm introduces a refinement phase subsequent to the local-moving phase, wherein vertices within each community undergo constrained merges to other sub-communities within their community bounds (obtained from the local-moving phase), starting from a singleton sub-community. This is performed in a randomized manner, with the probability of joining a neighboring sub-community within its community bound being proportional to the delta-modularity of the move. This facilitates the identification of sub-communities within those obtained from the local-moving phase.Once communities have converged, it is guaranteed that all vertices are optimally assigned, and all communities are subset optimal (Traag, Waltman, and van Eck 2019).
      Model       MSE       MAE       RMSE       p-value       R-squared
      GRU       0.8311       0.1264       0.9116       1.35e-18       0.3158
      RNN       0.9034       0.1190       0.9504       4.50e-24       0.3160
      LSTM       0.7921       0.1053       0.9736       1.19e-29       0.3254
      Ours       0.6021       0.0832       0.7923       2.19e-44       0.3775
Table 1: P/B ratio prediction performance on the benchmark datasets.
Mask Matrix Regularization
Recall the origin Lasso (Tibshirani 1996) and adaptive Lasso (Zou 2006) formulations:
@@
where $\mathbf{r}$ is the element-wise reciprocal of the pre-estimator of $\rho$.
After we extract the sparse connectivity structures $\mathcal{E}$ from $G$ using Leiden algorithm , We define the mask matrices formulation from (Ibrahim et al. 2022) as follows:
@@
These masking matrices $M$ are used to apply modified regularization penalties to $\rho$. As the masking entries increase, the corresponding penalties applied to the entries in $\rho$ also increase. Larger entries effectively zero out the corresponding elements in $\rho$, thereby reducing model complexity. With these additional masking weights, we use the newly redefined masked matrix regularizer from (Ibrahim et al. 2022) as follows:
@@
Based on the experiment results in Figure 2, our new masked regularizer were found to be effective in improving model performance by enforcing sparsity and preserving important structural information in the graph. This regularizer not only enhanced the interpretability of the learned representations but also demonstrated robustness against overfitting, particularly in high-dimensional and complex datasets.
Experiment
Datasets
@@
Our dataset is composed of two parts: BBT-FinCorpus (Lu et al. 2023) and Wanjuan1.0 (He et al. 2023). BBT-FinCorpus is a large-scale Chinese financial corpus that includes approximately 300GB of raw text data gathered from four distinct sources: (1) financial, political, and economic news published by mainstream media platforms over the past 20 years; (2) all announcements and financial reports from publicly listed companies; and (3) millions of research reports from financial institutes and consulting agencies. These sources provide a comprehensive dataset that covers a wide range of financial and economic topics, offering rich background information and multidimensional data for our research. The second part of the dataset comes from Wanjuan1.0, which includes 7 million news articles covering various fields across the Chinese internet. The combination of these two datasets gives us a broad and in-depth perspective, enabling a multi-faceted analysis and prediction of startup success.
Data Processing
@@
However, processing such a massive dataset using the GPT-4o mini API (OpenAI 2024) is quite expensive with costs amounting to \$5.00 per 1 million tokens for input and \$15.00 per 1 million tokens for output. To optimize processing efficiency and control costs, we implemented a series of data filtering measures. First, we filtered the data based on length, removing any text longer than 10,000 characters. This significantly reduced the data volume, making subsequent processing more manageable. Next, we focused on filtering data for over 300 companies that were listed on the Chinese A-share market between 2023 and 2024. We created a list of these company names and used it to search for matching fields in the news articles, removing any data that was entirely unrelated to these companies, thereby further streamlining the dataset.
@@
To ensure the relevance and quality of the data, we employed the Deepseek-chat (DeepSeek-AI et al. 2024) API’s JSON formatting output function to assign a relevance score ranging from 0 to 5 for each news article based on its importance relative to the company names. We chose to retain only those articles with higher relevance scores and removed those with a score of 0. This meticulous filtering process resulted in a refined dataset of 40,000 news articles, with an average length of 3,000 tokens per article.
@@
As shown in Table 1, our proposed GraphRAG augmented sequence-to-sequence LSTM method outperforms all baseline methods on this task. Specifically, we observed an approximately 16% improvement in R-squared compared to the best-performing LSTM model among the baselines, which fully demonstrates the effectiveness of our approach. This result highlights our model’s superior ability to capture complex inter-company relationships and handle sparse data, leading to significantly better prediction accuracy. Moreover, our model exhibited robustness and reliability across various data sparsity conditions, further validating the potential of GraphRAG technology in the venture capital domain.
Limitation
Data dependency is a major limiting factor. Although GraphRAG can effectively integrate multi-source information, its performance heavily depends on the quality and breadth of the data used. The costs associated with acquiring and processing unstructured text data are high, particularly during the data processing stage, where the expenses for large-scale text data processing are substantial. The indexing phase of GraphRAG introduces computational cost challenges due to its complexity and heavy computational requirements, which may limit its practical application in certain scenarios. Additionally, the knowledge graphs and other data generated without effective filtering can be large, leading to a significant increase in memory and computational resource consumption. The model’s generalization ability may also be affected by the dataset. While our model performs well on the dataset of Chinese A-share listed companies, its generalization ability across other datasets or market conditions requires further validation.
Future Work
While our proposed GraphRAG-enhanced multivariate time series analysis method has shown promise in predicting startup success, there are several areas that warrant further investigation to address current limitations and enhance the model’s capabilities.
Generalization Across Datasets: A key limitation of our current work is the reliance on non-structured data, which can be difficult to obtain consistently across different regions and industries. The dataset used in this study is primarily focused on the Chinese A-share market, which may limit the generalizability of our findings. Future research should focus on testing the model’s performance on datasets from other countries and markets to evaluate its robustness and adaptability across different economic and cultural environments.
Exploration of Alternative Soft Masking Techniques: The current approach utilizes a specific method for constructing the mask matrix based on the Leiden algorithm. However, there is potential to explore other soft masking techniques that might offer improved performance. Future work could investigate the use of different clustering algorithms to generate mask matrices, potentially enhancing the model’s ability to capture complex inter-company relationships.
Exploration of Advanced Graph Neural Networks: While GraphRAG has shown promise in enhancing time series predictions, there is potential to explore more advanced graph neural network architectures. For instance, Graph Attention Networks (GATs) or Graph Convolutional Networks (GCNs) could be integrated into the model to better capture complex inter-company relationships and improve predictive performance.
Conclusion
@@
In this paper, we proposed a seq2seq multivariate time series analysis method based on GraphRAG and LSTM to predict the success of startups in venture capital. By integrating GraphRAG with traditional time series models, we addressed the key issue often overlooked in previous studies-the relationships between companies, and the experimental results shows significant improvements across various evaluation metrics comparing to existing deep learning methods, indicating that GraphRAG possesses strong capabilities in capturing the complex dynamics of startup success. The core advantage of GraphRAG lies in its ability to enhance the model’s reasoning capabilities through the construction of knowledge graphs and information retrieval, enabling a better understanding of the complex competitive and cooperative relationships among startups and allows for a substantial improvement in prediction accuracy and robustness under conditions of sparse data. Our approach not only mitigates some challenges related to data sparsity but also provides more accurate decision support for venture capitalists.

### v4

@@ 3.2 Graph Clustering
The algorithm introduces a refinement phase subsequent to the local-moving phase, wherein vertices within each community undergo constrained merges to other sub-communities within their community bounds (obtained from the local-moving phase), starting from a singleton sub-community. This is performed in a randomized manner, with the probability of joining a neighboring sub-community within its community bound being proportional to the delta-modularity of the move. This facilitates the identification of sub-communities within those obtained from the local-moving phase. Once communities have converged, it is guaranteed that all vertices are optimally assigned, and all communities are subset optimal Traag et al. (2019).
Since startup success prediction in venture capital often uses a quarter as the time unit, the performance differences between time series models are usually minimal. Instead, we focus more on how to leverage information from the GraphRAG-produced knowledge graph to enhance the performance of time series models.
3.3 Mask Matrix Regularization
Recall the origin Lasso Tibshirani (1996) and adaptive Lasso Zou (2006) formulations:
@@ 3.3 Mask Matrix Regularization
where $\mathbf{r}$ is the element-wise reciprocal of the pre-estimator of $\rho$. After we extract the sparse connectivity structures $\mathcal{E}$ from $G$ using Leiden algorithm, we define the mask matrices formulation from Ibrahim et al. (2022) as follows:
@@
These masking matrices $M$ are used to apply modified regularization penalties to $\rho$. As the masking entries increase, the corresponding penalties applied to the entries in $\rho$ also increase. Larger entries effectively zero out the corresponding elements in $\rho$, thereby reducing model complexity. With these additional masking weights, we use the newly redefined masked matrix regularizer from Ibrahim et al. (2022) as follows:
@@
Based on the ablation study results in Section 4.4, our new masked regularizer were found to be effective in improving model performance by enforcing sparsity and preserving important structural information in the graph. This regularizer not only enhanced the interpretability of the learned representations but also demonstrated robustness against overfitting, particularly in high-dimensional and complex datasets.
      Model       MSE       MAE       RMSE       p-value       R-squared
      GRU       0.8311       0.1264       0.9116       1.35e-18       0.3158
      RNN       0.9034       0.1190       0.9504       4.50e-24       0.2860
      BERT       0.8054       0.1083       0.9467       1.22e-25       0.3095
      Transformer       0.7921       0.1021       0.9514       2.16e-24       0.3275
      LSTM       0.7687       0.1053       0.9736       1.19e-29       0.3354
      Ours       0.6021       0.0832       0.7923       2.19e-44       0.4075
Table 1: P/B ratio prediction performance on the benchmark datasets.
4 Experiment
4.1 Datasets
@@
Our dataset is composed of two parts: BBT-FinCorpus Lu et al. (2023) and Wanjuan1.0 He et al. (2023). BBT-FinCorpus is a large-scale Chinese financial corpus that includes approximately 300GB of raw text data gathered from four distinct sources: (1) financial, political, and economic news published by mainstream media platforms over the past 20 years; (2) all announcements and financial reports from publicly listed companies; and (3) millions of research reports from financial institutes and consulting agencies. These sources provide a comprehensive dataset that covers a wide range of financial and economic topics, offering rich background information and multidimensional data for our research. The second part of the dataset comes from Wanjuan1.0, which includes 7 million news articles covering various fields across the Chinese internet. The combination of these two datasets gives us a broad and in-depth perspective, enabling a multi-faceted analysis and prediction of startup success.
4.2 Data Processing Pipeline
@@
However, processing such a massive dataset using the GPT-4o mini API OpenAI (2024) is quite expensive with costs amounting to \$5.00 per 1 million tokens for input and \$15.00 per 1 million tokens for output. To optimize processing efficiency and control costs, we implemented a series of data filtering measures. First, we filtered the data based on length, removing any text longer than 10,000 characters. This significantly reduced the data volume, making subsequent processing more manageable. Next, we focused on filtering data for over 300 companies that were listed on the Chinese A-share market between 2023 and 2024. We created a list of these company names and used it to search for matching fields in the news articles, removing any data that was entirely unrelated to these companies, thereby further streamlining the dataset.
@@
To ensure the relevance and quality of the data, we employed the Deepseek-chat DeepSeek-AI et al. (2024) API’s JSON formatting output function to assign a relevance score ranging from 0 to 5 for each news article based on its importance relative to the company names. We chose to retain only those articles with higher relevance scores and removed those with a score of 0. This meticulous filtering process resulted in a refined dataset of 40,000 news articles, with an average length of 3,000 tokens per article.
@@ 4.3 Experimental Setup
As shown in Table 1, our proposed GraphRAG augmented sequence-to-sequence LSTM method outperforms all baseline methods on this task. Specifically, we observed an approximately 16% improvement in R-squared compared to the best-performing model among the baselines, which fully demonstrates the effectiveness of our approach. This result highlights our model’s superior ability to capture complex inter-company relationships and handle sparse data, leading to significantly better prediction accuracy. Moreover, our model exhibited robustness and reliability across various data sparsity conditions, further validating the potential of GraphRAG technology in the venture capital domain.
4.4 Ablation Study
We next demonstrate through two sets of ablation experiments that the information produced by the GraphRAG-generated knowledge graph effectively enhances the time series prediction of startup success. In the first set of experiments, we adjust the regularization strength of the Mask Matrix Regularization by gradually increasing the size of the regularization term, thereby controlling the amount of knowledge graph information introduced into the time series model. In the second set of experiments, we randomly remove edges from the knowledge graph, representing relationships such as collaboration, competition, and investment between entities, to show that incorporating these relationships significantly improves the performance of time series prediction for startup success.
$\lambda$ MSE MAE R-squared
0.1 0.6213 0.0901 0.3552
0.5 0.6135 0.0876 0.3948
1.0 0.6054 0.0854 0.4019
5.0 0.6027 0.0840 0.4061
10.0 0.6021 0.0832 0.4075
Table 2: Effect of Regularization Strength $\lambda$ (optimal 10) on MSE, MAE, and R-squared
$\alpha$ MSE MAE R-squared
0.2 0.6589 0.0987 0.3402
0.4 0.6348 0.0934 0.3625
0.6 0.6175 0.0898 0.3803
0.8 0.6063 0.0865 0.3971
1.0 0.6021 0.0832 0.4075
Table 3: Effect of Proportion of Unremoved Nodes $\alpha$ on MSE, MAE, and R-squared
The ablation studies confirm that integrating knowledge graph produced by GraphRAG significantly improves startup success prediction. Adjusting regularization strength ($\lambda$) enhances accuracy, while removing key graph edges reduces performance, highlighting the importance of relational data like collaboration and competition. These results validate the effectiveness of our GraphRAG-based approach in handling sparse data and low signal challenges in venture capital analysis.
5 Limitation
GraphRAG’s performance relies heavily on data quality, with high costs for processing unstructured text and computational challenges during indexing. Large, unfiltered knowledge graphs increase resource consumption, and the model’s generalization beyond Chinese A-share datasets needs further validation.
6 Conclusion
@@
This paper proposed a GraphRAG augmented Seq2Seq multivariate time series method to predict startup success in venture capital. By integrating GraphRAG with traditional time series models, we tackled a key issue often overlooked in previous studies: the inter-company relationships. Experimental results show significant improvements over existing methods, with GraphRAG enhancing reasoning by leveraging knowledge graphs and retrieval to capture competitive and cooperative dynamics, improving accuracy and robustness under sparse data conditions. Our approach not only mitigates some challenges related to data sparsity but also provides more accurate decision support for venture capitalists.